# Implicit Bias of Per-sample Adam on Separable Data: Departure from the Full-batch Regime

### v2

@@ Abstract
Adam (Kingma and Ba, 2015) is the de facto optimizer in deep learning, yet its theoretical understanding remains limited. Prior analyses show that Adam favors solutions aligned with $\ell_{\infty}$-geometry, but these results are restricted to the full-batch regime. In this work, we study the implicit bias of incremental Adam (using one sample per step) for logistic regression on linearly separable data, and we show that its bias can deviate from the full-batch behavior. To illustrate this, we construct a class of structured datasets where incremental Adam provably converges to the $\ell_{2}$-max-margin classifier, in contrast to the $\ell_{\infty}$-max-margin bias of full-batch Adam. For general datasets, we develop a proxy algorithm that captures the limiting behavior of incremental Adam as $\beta_{2}\to 1$ and we characterize its convergence direction via a data-dependent dual fixed-point formulation. Finally, we prove that, unlike Adam, Signum (Bernstein et al., 2018) converges to the $\ell_{\infty}$-max-margin classifier for any batch size by taking $\beta$ close enough to 1. Overall, our results highlight that the implicit bias of Adam crucially depends on both the batching scheme and the dataset, while Signum remains invariant.
@@ 1 Introduction
A canonical setting for studying implicit bias is linear classification on separable data with logistic loss. In this setup, achieving zero training loss requires the model’s weights to diverge to infinity, making the direction of convergence—which defines the decision boundary—the key object of study. Seminal work by Soudry et al. (2018) establishes that gradient descent (GD) converges to the $\ell_{2}$-max-margin solution. This foundational result has inspired extensive research extending the analysis to neural networks, alternative optimizers, and other loss functions (Gunasekar et al., 2018b; Ji and Telgarsky, 2019, 2020; Lyu and Li, 2020; Chizat and Bach, 2020; Yun et al., 2021). In this work, we revisit the simplest setting—linear classification on separable data—to examine how the choice of optimizer shapes implicit bias.
Among modern optimization algorithms, Adam (Kingma and Ba, 2015) is one of the most widely used, making its implicit bias particularly important to understand. Zhang et al. (2024a) show that, unlike GD, full-batch Adam converges in direction to the $\ell_{\infty}$-max-margin solution. This behavior is closely related to sign gradient descent (SignGD), which can be interpreted as normalized steepest descent with respect to the $\ell_{\infty}$-norm and is also known to converge to the $\ell_{\infty}$-max-margin direction (Gunasekar et al., 2018a; Fan et al., 2025). Xie et al. (2025) further attribute Adam’s empirical success in language model training to its ability to exploit the favorable $\ell_{\infty}$-geometry of the loss landscape.
Yet, prior work on implicit bias in linear classification has almost exclusively focused on the full-batch setting. In contrast, modern training relies on stochastic mini-batches, a regime where theoretical understanding remains limited. Notably, Nacson et al. (2019) show that SGD preserves the same $\ell_{2}$-max-margin bias as GD, suggesting that mini-batching may not alter an optimizer’s implicit bias. But does this extend to adaptive methods such as Adam?
@@
Perhaps surprisingly, we find that the answer is no. Our experiments (Figure 1) illustrate that when trained on Gaussian data, full-batch Adam converges to the $\ell_{\infty}$-max-margin direction, whereas mini-batch Adam variants with batch size $1$ converge closer to the $\ell_{2}$-max-margin direction. To explain this phenomenon, we develop a theoretical framework for analyzing the implicit bias of mini-batch Adam, focusing on the batch size $1$ case as a representative contrast to the full-batch regime. To the best of our knowledge, this work provides the first theoretical evidence that Adam’s implicit bias is fundamentally altered in the mini-batch setting.
@@
  • We demonstrate a sharp contrast between full-batch and mini-batch Adam using a family of structured datasets, Generalized Rademacher (GR) data. On GR data, we prove that incremental Adam converges to the $\ell_{2}$-max-margin solution, while full-batch Adam converges to the $\ell_{\infty}$-max-margin solution (see Section 3).
  • For general datasets, we introduce a uniform-averaging proxy that predicts the limiting behavior of incremental Adam as $\beta_{2}\to 1$. We characterize its convergence direction as the primal solution of an optimization problem defined by a dual fixed-point equation (see Section 4).
@@
Figure 1: Mini-batch Adam loses the $\ell_{\infty}$-max-margin bias of full-batch Adam. Cosine similarity between the weight vector and the $\ell_{2}$-max-margin (left) and $\ell_{\infty}$-max-margin (right) solutions in a linear classification task on $10$ data points drawn from the $50$-dimensional standard Gaussian. Full-batch Adam with $(\beta_{1},\beta_{2})=(0.9,0.95)$ converges to the $\ell_{\infty}$-max-margin solution, whereas mini-batch variants with batch size $1$ converge closer to the $\ell_{2}$-max-margin direction. See Appendix C for experimental details.
@@ 2 How Can We Approximate Without-replacement Adam?
For a vector $\mathbf{v}$}, let ${\mathbf{v}}[k]$ denote its $k$-th entry, ${\mathbf{v}}_{t}$ its value at time step $t$, and ${\mathbf{v}}_{r}^{s}\triangleq{\mathbf{v}}_{rN+s}$ unless stated otherwise. For a matrix $\mathbf{M}$}, let ${\mathbf{M}}[i,j]$ denote its $(i,j)$-th entry. We use $\Delta^{N-1}$ to denote the probability simplex in ${\mathbb{R}}^{N}$. Let $N]=\{0,1,\cdots,N-1\}$ denote the set of the first $N$ non-negative integers. For a PSD matrix $\mathbf{M}$}, define the energy norm as $\|{\mathbf{x}}\|_{\mathbf{M}}\triangleq\sqrt{{\mathbf{x}}^{\top}{\mathbf{M}}{\mathbf{x}}}$. For vectors, $\sqrt{\cdot}$, $(\cdot)^{2}$, and $\frac{\cdot}{\cdot}$ operations are applied entry-wise unless stated otherwise. Given two functions $f(t),g(t)$, we denote $f(t)=\mathcal{O}(g(t))$ if there exist $C,T>0$ such that $t\geq T$ implies $|f(t)|\leq C|g(t)|$. For two vectors $\mathbf{v}$} and $\mathbf{w}$}, we denote ${\mathbf{v}}\propto{\mathbf{w}}$ if ${\mathbf{v}}=c\cdot{\mathbf{w}}$ for a positive scalar $c>0$. Let $r=a\bmod b$ with $0\leq r<b$ denote the unique integer remainder when dividing $a$ by $b$.
@@
We focus on incremental Adam (Inc-Adam), which processes mini-batch gradients sequentially from indices 0 to N1N-1 in each epoch. Studying Inc-Adam provides a tractable way to understand the implicit bias of mini-batch Adam: our experiments show that its iterates converge in directions closely aligned with mini-batch Adam of batch size $1$ under both with-replacement and random-reshuffling sampling. Sharing the same mini-batch accumulation mechanism, Inc-Adam serves as a faithful surrogate for theoretical analysis. Pseudocodes for Inc-Adam and full-batch deterministic Adam (Det-Adam) are given in Algorithms 1 and 2.
@@
We primarily focus on binary linear classification tasks. To be specific, training data are given by $\{({\mathbf{x}}_{i},y_{i})\}_{i\in[N]}$, where ${\mathbf{x}}_{i}\in{\mathbb{R}}^{d}$, $y_{i}\in\{-1,+1\}$. We aim to find a linear classifier $\mathbf{w}$} which minimizes the loss
@@
where $\ell:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is a surrogate loss for classification accuracy and $\mathcal{L}_{i}({\mathbf{w}})=\ell(y_{i}\langle{\mathbf{w}},{\mathbf{x}}_{i}\rangle)$ denotes the loss value on the $i$-th data point. Without loss of generality, we assume $y_{i}=+1$, since we can newly define $\tilde{{\mathbf{x}}}_{i}=y_{i}{\mathbf{x}}_{i}$. In this paper, we consider two loss functions $\ell\in\{\ell_{\text{exp}},\ell_{\text{log}}\}$, where $\ell_{\text{exp}}(z)=\exp(-z)$ denotes the exponential loss and $\ell_{\text{log}}(z)=\log(1+e^{-z})$ denotes the logistic loss.
@@
Assumption 2.2 (Nonzero entries).
${\mathbf{x}}_{i}[k]\neq 0$ for all $i\in[N],k\in[d]$.
@@
The dependence of the Adam update on the full gradient history makes its asymptotic analysis largely intractable. We address this challenge with the following propositions, which show that the epoch-wise updates of Inc-Adam and the updates of Det-Adam can be approximated by a function that depends only on the current iterate. This result forms a cornerstone of our future analysis.
@@
Propositions 2.4 and 2.5 reveal a fundamental discrepancy between the behavior of Det-Adam and that of Inc-Adam. Proposition 2.4 demonstrates that Det-Adam can be approximated by SignGD, which has been reported by previous works (Balles and Hennig, 2018; Zou et al., 2023). Note that the condition is not satisfied when $\nabla\mathcal{L}({\mathbf{w}}_{t})[k]$ decays at a rate on the order of $\eta_{t}^{1/2}\mathcal{L}({\mathbf{w}}_{t})$, which often calls for a more detailed analysis (see Zhang et al. (2024a, Lemma 6.2)). Such an analysis establishes that Det-Adam asymptotically finds an $\ell_{\infty}$-max-margin solution, a property that holds regardless of the choice of momentum hyperparameters satisfying $\beta_{1}\leq\beta_{2}$ (Zhang et al., 2024a).
@@ 3 Warmup: Structured Data
We define Generalized Rademacher (GR) data as a set of vectors $\{{\mathbf{x}}_{i}\}_{i\in[N]}$ which satisfy $|{\mathbf{x}}_{i}[k]|=|{\mathbf{x}}_{i}[l]|,\forall k,l\in[d]$, for each $i\in[N]$. We also assume that GR data satisfy Assumptions 2.2 and 2.1, unless otherwise specified.
Applying Proposition 2.5 to the GR dataset, we obtain the following corollary.
@@
Consider Inc-Adam iterates $\{{\mathbf{w}}_{t}\}_{t=0}^{\infty}$ on GR data. Then, under Assumptions 2.1, 2.2 and 2.3, the epoch-wise update ${\mathbf{w}}_{r+1}^{0}-{\mathbf{w}}_{r}^{0}$ can be approximated by weighted normalized GD, i.e.,
${\mathbf{w}}_{r+1}^{0}-{\mathbf{w}}_{r}^{0}=-\eta_{rN}\left(\frac{\sum_{i\in[N]}a_{i}(r)\nabla\mathcal{L}_{i}({\mathbf{w}}_{r}^{0})}{\|\nabla\mathcal{L}({\mathbf{w}}_{r}^{0})\|_{2}}+\bm{\epsilon}_{r}\right),$ (3)
@@
Although the using a structured dataset simplifies the denominator in Equation 2, the dynamics are still governed by weighted GD, which requires careful analysis. Prior work studies the implicit bias of weighted GD, particularly in the context of importance weighting (Xu et al., 2021; Zhai et al., 2023), but these analysis typically assume that the weights are constant or convergent. In our setting, the weight $a_{i}(r)$ varies with the epoch count $r$. We address this challenge and characterize the implicit bias of Inc-Adam on the GR data as follows.
@@
Consider Inc-Adam iterates $\{{\mathbf{w}}_{t}\}_{t=0}^{\infty}$ with $\beta_{1}\leq\beta_{2}$ on GR data under Assumptions 2.2, 2.1 and 2.3. If (a) $\mathcal{L}({\mathbf{w}}_{t})\rightarrow 0$ as $t\rightarrow\infty$ and (b) $\eta_{t}=(t+2)^{-a}$ for $a\in(2/3,1]$, then it satisfies
@@
where $\hat{{\mathbf{w}}}_{\ell_{2}}$ denotes the (unique) $\ell_{2}$-max-margin solution of GR data $\{{\mathbf{x}}_{i}\}_{i\in[N]}$.
The analysis in Theorem 3.3 relies on Corollary 3.2, which ensures that the weights $a_{i}(r)$ are bounded by two positive constants, $c_{1}$ and $c_{2}$. This condition is crucial to prevent any individual data from having a vanishing contribution, which could cause the Inc-Adam iterates to deviate from the $\ell_{2}$-max-margin direction. Furthermore, the controlled learning rate schedule is key to bounding the $\bm{\epsilon}_{r}$ term in our analysis. The proof and further discussion are deferred to Appendix E. As shown in Figure 2, our experiments on GR data confirm that mini-batch Adam with batch size $1$ converges in direction to the $\ell_{2}$-max-margin classifier, in contrast to the $\ell_{\infty}$-bias of full-batch Adam.
Notably, Theorem 3.3 holds for any choice of momentum hyperparameters satisfying $\beta_{1}\leq\beta_{2}$; see Figure 9 in Appendix B for empirical evidence. This invariance of the bias arises from the structure of GR data, which removes the coordinate adaptivity that momentum hyperparameters would normally affect. For general datasets, the invariance no longer holds; the adaptivity persists and varies with the choice of momentum hyperparameters, as discussed in Appendix A. In the next section, we introduce a proxy algorithm to study the regime where $\beta_{2}$ is close to $1$ and characterize its implicit bias.
Figure 2: Mini-batch Adam converges to the $\ell_{2}$-max-margin solution on the GR dataset. We train on the dataset ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{1}=(4,-4,4,-4)$. Variants of mini-batch Adam with batch size 11 consistently converge to the $\ell_{2}$-max-margin direction, while full-batch Adam converges to the $\ell_{\infty}$-max-margin direction.

### v3

@@ Abstract
Adam (Kingma and Ba, 2015) is the de facto optimizer in deep learning, yet its theoretical understanding remains limited. Prior analyses show that Adam favors solutions aligned with $\ell_{\infty}$-geometry, but these results are restricted to the full-batch regime. In this work, we study the implicit bias of incremental Adam (using one sample per step) for logistic regression on linearly separable data, and show that its bias can deviate from the full-batch behavior. As an extreme example, we construct datasets on which incremental Adam provably converges to the $\ell_{2}$-max-margin classifier, in contrast to the $\ell_{\infty}$-max-margin bias of full-batch Adam. For general datasets, we characterize its bias using a proxy algorithm for the $\beta_{2}\to 1$ limit. This proxy maximizes a data-adaptive Mahalanobis-norm margin, whose associated covariance matrix is determined by a data-dependent dual fixed-point formulation. We further present concrete datasets where this bias reduces to the standard $\ell_{2}$- and $\ell_{\infty}$-max-margin classifiers. As a counterpoint, we prove that Signum (Bernstein et al., 2018) converges to the $\ell_{\infty}$-max-margin classifier for any batch size. Overall, our results highlight that the implicit bias of Adam crucially depends on both the batching scheme and the dataset, while Signum remains invariant.
@@ 1 Introduction
A canonical setting for studying implicit bias is linear classification on separable data with logistic loss. In this setup, achieving zero training loss requires the model’s weights to diverge to infinity, making the directional convergence—which defines the decision boundary—the key object of study. Seminal work by Soudry et al. (2018) establishes that gradient descent (GD) directionally converges to the $\ell_{2}$-max-margin solution. This foundational result has inspired extensive research extending the analysis to neural networks, alternative optimizers, and other loss functions (Gunasekar et al., 2018a; Ji and Telgarsky, 2019; 2020; Lyu and Li, 2020; Chizat and Bach, 2020; Yun et al., 2021). In this work, we revisit the simplest setting—linear classification on separable data—to examine how the choice of optimizer shapes implicit bias.
Among modern optimization algorithms, Adam (Kingma and Ba, 2015) is one of the most widely used, making its implicit bias particularly important to understand. Zhang et al. (2024a) show that, unlike GD, full-batch Adam converges in direction to the $\ell_{\infty}$-max-margin solution. This behavior is closely related to sign gradient descent (SignGD), which can be interpreted as normalized steepest descent in the $\ell_{\infty}$-norm and is also known to converge to the $\ell_{\infty}$-max-margin direction (Gunasekar et al., 2018b; Fan et al., 2025). Xie et al. (2025) further attribute Adam’s empirical success in language model training to its ability to exploit the favorable $\ell_{\infty}$-geometry of the loss landscape.
Yet, prior work on implicit bias in linear classification has almost exclusively focused on the full-batch setting. In contrast, modern training relies on stochastic mini-batches, a regime where theoretical understanding remains limited. Notably, Nacson et al. (2019) show that SGD with an arbitrary batch size preserves the same $\ell_{2}$-max-margin bias as GD, suggesting that mini-batching may not alter an optimizer’s implicit bias. But does this extend to adaptive methods such as Adam?
@@
Perhaps surprisingly, we find that the answer is no. Our experiments (Figure 1) illustrate that when trained on Gaussian data, full-batch Adam converges to the $\ell_{\infty}$-max-margin direction, whereas Adam variants with a batch size of $1$ converge to a different direction, which is even closer to the $\ell_{2}$-max-margin solution. To explain this phenomenon, we develop a theoretical framework for analyzing the implicit bias of mini-batch Adam and focus on the batch size $1$ case as a representative contrast to the full-batch regime. To the best of our knowledge, this work provides the first theoretical evidence describing the data-dependent characterization of the implicit bias of Adam with a batch size of $1$.
Figure 1: Mini-batch Adam loses the $\ell_{\infty}$-max-margin bias of full-batch Adam. Cosine similarity between the weight vector and the $\ell_{2}$-max-margin (left) and $\ell_{\infty}$-max-margin (right) solutions in a linear classification task on $10$ data points drawn from the $50$-dimensional standard Gaussian. Full-batch Adam with $(\beta_{1},\beta_{2})=(0.9,0.95)$ converges to the $\ell_{\infty}$-max-margin solution, whereas mini-batch variants with a batch size of $1$ converge to a different direction (see Section 4 for the detailed characterization). See Appendix C for experimental details.
@@
  • We demonstrate a sharp contrast between full-batch and mini-batch Adam using a family of structured datasets, Scaled Rademacher (SR) data. On SR data, we prove that incremental Adam converges to the $\ell_{2}$-max-margin solution, while full-batch Adam converges to the $\ell_{\infty}$-max-margin solution (see Section 3).
  • For general dataset, we introduce a uniform-averaging proxy that characterizes the limiting behavior of incremental Adam as $\beta_{2}\to 1$. We identify its convergence direction as the solution of a data-adaptive margin-maximization problem, induced by a Mahalanobis norm whose covariance matrix is determined by a data-dependent dual fixed-point equation. We further present concrete datasets where this bias reduces to the standard $\ell_{2}$- and $\ell_{\infty}$-max-margin classifiers (see Section 4).
@@ 2 How Can We Approximate Without-replacement Adam?
For a vector $\mathbf{v}$}, let ${\mathbf{v}}[k]$ denote its $k$-th entry. For a matrix $\mathbf{M}$}, let ${\mathbf{M}}[i,j]$ denote its $(i,j)$-th entry. We use $\Delta^{N-1}$ to denote the probability simplex in ${\mathbb{R}}^{N}$. Let $N]=\{0,1,\cdots,N-1\}$ denote the set of the first $N$ non-negative integers. For a PSD matrix $\mathbf{M}$}, define the Mahalanobis norm as $\|{\mathbf{x}}\|_{\mathbf{M}}\triangleq\sqrt{{\mathbf{x}}^{\top}{\mathbf{M}}{\mathbf{x}}}$. For vectors, $\sqrt{\cdot}$, $(\cdot)^{2}$, and $\frac{\cdot}{\cdot}$ operations are applied entry-wise unless stated otherwise. Given two functions $f(t),g(t)$, we denote $f(t)=\mathcal{O}(g(t))$ if there exist $C,T>0$ such that $t\geq T$ implies $|f(t)|\leq C|g(t)|$. For two vectors $\mathbf{v}$} and $\mathbf{w}$}, we denote ${\mathbf{v}}\propto{\mathbf{w}}$ if ${\mathbf{v}}=c\cdot{\mathbf{w}}$ for a positive scalar $c>0$. Let $r=a\bmod b$ denote the remainder when dividing $a$ by $b$, i.e., $0\leq r<b$.
@@
We focus on incremental Adam (Inc-Adam), which processes mini-batch gradients sequentially from indices 0 to N1N-1 in each epoch. Studying Inc-Adam provides a tractable way to understand the implicit bias of mini-batch Adam under various sampling schemes: our experiments show that Inc-Adam’s directional convergence closely aligns with that of Adam with a batch size of $1$ under both sampling with replacement and random-reshuffling. Sharing the same mini-batch accumulation mechanism, Inc-Adam serves as a faithful surrogate for theoretical analysis. Pseudocodes for Inc-Adam and full-batch deterministic Adam (Det-Adam) are given in Algorithms 1 and 2.
@@
We primarily focus on binary linear classification tasks. To be specific, training data are given by $\{({\mathbf{x}}_{i},y_{i})\}_{i\in[N]}$, where ${\mathbf{x}}_{i}\in{\mathbb{R}}^{d}$, $y_{i}\in\{-1,+1\}$, and $N$ is the number of data points. We aim to find a linear classifier $\mathbf{w}$} which minimizes the loss
@@
where $\ell:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is a surrogate loss for classification accuracy and $\mathcal{L}_{i}({\mathbf{w}})=\ell(y_{i}\langle{\mathbf{w}},{\mathbf{x}}_{i}\rangle)$ denotes the loss value on the $i$-th data point. Without loss of generality, we assume $y_{i}=+1$, since we can newly define $\tilde{{\mathbf{x}}}_{i}=y_{i}{\mathbf{x}}_{i}$. In this paper, we consider two loss functions $\ell\in\{\ell_{\text{exp}},\ell_{\text{log}}\}$, where $\ell_{\text{exp}}(z)=\exp(-z)$ denotes the exponential loss and $\ell_{\text{log}}(z)=\log(1+e^{-z})$ denotes the logistic loss. Given a sequence of vectors $\{{\mathbf{v}}_{t}\}_{t=0}^{\infty}$, for notational simplicity, let ${\mathbf{v}}_{r}^{s}\triangleq{\mathbf{v}}_{rN+s}$ denote the $s$-th element in the $r$-th group of $N$ consecutive terms.
@@
Assumption 2.2 (Nonzero coordinates).
For all $i\in[N]$ and $k\in[d]$, ${\mathbf{x}}_{i}[k]\neq 0$.
@@
The dependence of the Adam update on the full gradient history makes its asymptotic analysis largely intractable. We address this challenge with the following propositions, which show that the epoch-wise updates of Inc-Adam and the updates of Det-Adam can be approximated by a function that depends only on the current iterate. Detailed proofs are deferred to Appendix D.
@@
Propositions 2.4 and 2.5 reveal a fundamental discrepancy between the behavior of Det-Adam and one of Inc-Adam. Proposition 2.4 demonstrates that Det-Adam can be approximated by SignGD, which has been reported by previous works (Balles and Hennig, 2018; Zou et al., 2023). Note that the condition is not satisfied when $\nabla\mathcal{L}({\mathbf{w}}_{t})[k]$ decays at a rate on the order of $\eta_{t}^{1/2}\mathcal{L}({\mathbf{w}}_{t})$, which often calls for a more detailed analysis (see Zhang et al. (2024a, Lemma 6.2)). Such an analysis establishes that Det-Adam asymptotically finds an $\ell_{\infty}$-max-margin solution, a property that holds regardless of the choice of momentum hyperparameters satisfying $\beta_{1}\leq\beta_{2}$ (Zhang et al., 2024a).
@@ 3 Warmup: Structured Data
We define Scaled Rademacher (SR) data as a set of vectors $\{{\mathbf{x}}_{i}\}_{i\in[N]}$ which satisfy $|{\mathbf{x}}_{i}[k]|=|{\mathbf{x}}_{i}[l]|,\forall k,l\in[d]$, for each $i\in[N]$. We also assume that SR data satisfy Assumptions 2.2 and 2.1, unless otherwise specified.
Applying Proposition 2.5 to the SR dataset, we obtain the following corollary.
@@
Consider Inc-Adam iterates $\{{\mathbf{w}}_{t}\}_{t=0}^{\infty}$ on SR data. Then, under Assumptions 2.2 and 2.3, the epoch-wise update ${\mathbf{w}}_{r+1}^{0}-{\mathbf{w}}_{r}^{0}$ can be approximated by weighted normalized GD, i.e.,
${\mathbf{w}}_{r+1}^{0}-{\mathbf{w}}_{r}^{0}=-\eta_{rN}\left(\sum_{i\in[N]}\frac{a_{i}(r)}{\|\nabla\mathcal{L}({\mathbf{w}}_{r}^{0})\|_{2}}\nabla\mathcal{L}_{i}({\mathbf{w}}_{r}^{0})+\bm{\epsilon}_{r}\right),$ (3)
@@
Although using a structured dataset simplifies the denominator in Equation 2, the dynamics are still governed by weighted GD, which requires a careful analysis. Prior work studies the implicit bias of weighted GD, particularly in the context of importance weighting (Xu et al., 2021; Zhai et al., 2023), but these analysis typically assume that the weights are constant or convergent. In our setting, the weight $a_{i}(r)$ varies with the epoch count $r$. We address this challenge and characterize the implicit bias of Inc-Adam on the SR data as follows.
@@
Consider Inc-Adam iterates $\{{\mathbf{w}}_{t}\}_{t=0}^{\infty}$ with $\beta_{1}\leq\beta_{2}$ on SR data under Assumptions 2.2, 2.1 and 2.3. If (a) $\mathcal{L}({\mathbf{w}}_{t})\rightarrow 0$ as $t\rightarrow\infty$ and (b) $\eta_{t}=(t+2)^{-a}$ for $a\in(2/3,1]$, then it satisfies
@@
where $\hat{{\mathbf{w}}}_{\ell_{2}}$ denotes the (unique) $\ell_{2}$-max-margin solution of SR data $\{{\mathbf{x}}_{i}\}_{i\in[N]}$.
The analysis in Theorem 3.3 relies on Corollary 3.2, which ensures that the weights $a_{i}(r)$ are bounded by two positive constants, $c_{1}$ and $c_{2}$. This condition is crucial to prevent any individual data from having a vanishing contribution, which could cause the Inc-Adam iterates to deviate from the $\ell_{2}$-max-margin direction. Furthermore, the controlled learning rate schedule is key to bounding the $\bm{\epsilon}_{r}$ term in our analysis. The proof and further discussion are deferred to Appendix E. As shown in Figure 2, our experiments on SR data confirm that mini-batch Adam with a batch size of $1$ converges in direction to the $\ell_{2}$-max-margin classifier, in contrast to the $\ell_{\infty}$-bias of full-batch Adam.
Notably, Theorem 3.3 holds for any choice of momentum hyperparameters satisfying $\beta_{1}\leq\beta_{2}$; see Figure 9 in Appendix B for empirical evidence. This invariance of the bias arises from the structure of SR data, which removes the coordinate adaptivity that momentum hyperparameters would normally affect. For general datasets, the invariance no longer holds; the adaptivity persists and varies with the choice of momentum hyperparameters, as discussed in Appendix A. In the next section, we introduce a proxy algorithm to study the regime where $\beta_{2}$ is close to $1$ and characterize its implicit bias.
Figure 2: Mini-batch Adam converges to the $\ell_{2}$-max-margin solution on the SR dataset. We train on the dataset ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{1}=(4,-4,4,-4)$. Variants of mini-batch Adam with a batch size of 11 consistently converge to the $\ell_{2}$-max-margin direction, while full-batch Adam converges to the $\ell_{\infty}$-max-margin direction.